It takes 2.60 s for a small ball with a mass of 0.060 kg released from rest from a tall building to reach the ground. Calculate the height from which the ball is released. If that ball had been released from the same height, but this time from the surface of the moon, how long would it have taken for the ball to hit the ground?


1
Expert's answer
2019-09-24T09:02:02-0400

For the Earth (assuming no air friction)


"h=\frac{gt^2}{2}=\frac{9.8\cdot2.6^2}{2}=33.124 m"


For the Moon


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\cdot33.124}{1.62}}=6.39 s"


Answer. 33.124 m; 6.39s